# The synthetic instrument: From sparse association to sparse causation

### v2

@@ Abstract
In many observational studies, researchers are often interested in studying the effects of multiple exposures on a single outcome. Standard approaches for high-dimensional data, such as the lasso, assume that the associations between the exposures and the outcome are sparse. However, these methods do not estimate causal effects in the presence of unmeasured confounding. In this paper, we consider an alternative approach that assumes the causal effects under consideration are sparse. We show that with sparse causation, causal effects are identifiable even with unmeasured confounding. At the core of our proposal is a novel device called the synthetic instrument, which, in contrast to standard instrumental variables, can be constructed using the observed exposures directly. We demonstrate that under the assumption of sparse causation, the problem of causal effect estimation can be formulated as an $\ell_{0}$-penalization problem and can thus be solved efficiently using off-the-shelf software. Simulations show that our approach outperforms state-of-the-art methods in both low-dimensional and high-dimensional settings. We further illustrate our method using a mouse obesity dataset.
@@
1Department of Statistics and Data Science, Washington University in St. Louis, St. Louis, Missouri, U.S.A.
@@ 1 Introduction
Sparsity is a common assumption in modern statistical learning literature, as it allows for variable selection in underlying models and enhances the interpretability of parameter estimates. For example, the lasso method (Tibshirani, 1996) assumes that the underlying associations between a single outcome and potentially high-dimensional predictors are sparse. In other words, only a small number of predictors have non-zero associations with the outcome. Hastie et al. (2009) summarize the philosophy behind these methods as the “bet on sparsity” principle: Use a procedure that performs well in sparse problems, as no procedure performs well in dense problems. Methods like the lasso perform well when the associations between the outcome and predictors are sparse and have gained significant popularity in recent decades.
The “bet on sparsity” principle does not restrict the type of “sparse problems” that can be considered. Instead of focusing solely on sparse associations, there is a growing body of literature that emphasizes sparse causation, where only a fraction of the exposures have non-zero causal effects on the outcomes (e.g., Spirtes and Glymour, 1991; Claassen et al., 2013; Wang et al., 2017; Miao et al., 2023; Zhou et al., 2024). This assumption is arguably more interpretable and plausible in many data applications. For example, suppose one is interested in the relationships between the expression of multiple genes and a phenotype of interest, such as lung cancer. Current biological knowledge suggests that not all genes in the human genome may affect lung cancer (e.g., Kanwal et al., 2017). However, this does not imply the assumption of sparse association, as unmeasured confounding often leads to spurious correlations between the exposures and the outcome.
For example, consider a linear structural model (Pearl, 2013) with a $p$-dimensional exposure $X=(X_{1},\ldots,X_{p})^{\mathrm{\scriptscriptstyle T}}$, an outcome $Y$, and a $q$-dimensional latent variable $U$:
@@
where $\Lambda\in\mathbb{R}^{p\times q}$, $\beta\in\mathbb{R}^{p}$, and $\gamma\in\mathbb{R}^{q}$ are coefficients, and $\epsilon_{x}=(\epsilon_{1},\ldots,\epsilon_{p})^{\mathrm{\scriptscriptstyle T}}$, $\epsilon_{y}$, and $U$ are uncorrelated with each other. Under this model, the spurious correlations due to unmeasured confounding, characterized by $\text{Cov}(X)^{-1}\Lambda\gamma$, are typically dense. As a result, the association between $X$ and $Y$ is not sparse, even if the causal effect $\beta$ is sparse.
Identification and estimation of the causal parameter $\beta$ are non-trivial due to the presence of unmeasured confounding by $U$. Our contributions in this paper are twofold. First, under an additional plurality condition, we establish that the parameter $\beta$ in model (3) is identifiable if and only if $||\beta||_{0}<p-q$. Our assumption on the sparsity level is both necessary and sufficient, representing a significant improvement over assumptions previously introduced in the literature; see Section 1.1 for details. Remarkably, in contrast to many other identification assumptions in causal inference, it is possible to consistently test the sparsity assumption from data.
Second, we develop a synthetic two-stage regularized regression approach for estimating $\beta$, with a first-stage ordinary least squares and a second-stage $\ell_{0}$-penalized regression. The key technique behind our results is a novel device we call the synthetic instrument. Unlike standard instrumental variables, the synthetic instrument is constructed from a subset of exposures, allowing us to identify the causal effects without resorting to exogenous variables. Our procedure comes with lasso-type theoretical guarantees in both low- and high-dimensional settings.
@@ 1.1 Related works
Our proposal is related to recent work dealing with multivariate hidden confounding. Ćevid et al. (2020) and Guo et al. (2022) propose a spectral deconfounding method for estimating $\beta$ under a high-dimensional model. Their method assumes a dense confounding model, which is only possible under a high-dimensional regime where $p$ tends to infinity with the sample size and the magnitude of spurious associations tends to zero. Bing et al. (2022) consider a more general setup than ours, in which they also allow the outcome to be multivariate. However, they aim to identify the projection of $\beta$ onto a related space rather than the causal parameter $\beta$ itself. Chandrasekaran et al. (2010) study a related problem under the assumption that $(X,Y)$ are normally distributed. Under this assumption, they not only identify the effect of $X$ on $Y$ but also recover the covariance among components of $X$ conditional on $U$.
@@
The estimation problem of $\beta$ can be framed within the context of causal inference with unmeasured confounding. Currently, the most popular approach in practice is the instrumental variable (IV) framework, which uses information from an exogenous variable known as an IV to identify causal effects. (e.g., Angrist et al., 1996; Wang and Tchetgen Tchetgen, 2018). Another approach that has gained attention recently is the proximal causal inference framework (Tchetgen Tchetgen et al., 2024), which uses information from ancillary variables known as negative control exposures and outcomes to remove bias due to unmeasured confounding. Compared with these frameworks, our approach does not rely on the collection of additional ancillary variables, which can be challenging in many practical settings. Instead, we rely on the availability of multiple exposures and the sparsity assumption for identification and estimation.
@@
Recently, there has been a strand of literature that seeks to identify the causal effects of multiple exposures. Wang and Blei (2019) popularized this setting by proposing the so-called deconfounder method, which first obtains an estimate $\widehat{U}$ for the unmeasured confounder and then adjusts for $\widehat{U}$ using standard regression methods. However, it has been pointed out that under this setting, without further assumptions, the causal effect $\beta$ is not identifiable (D’Amour, 2019; Ogburn et al., 2020). Kong et al. (2022) show that under model (1) and a binary choice model for the outcome with a non-probit link, the causal effects are identifiable. Their identification results, however, apply only to binary outcomes and do not lead to straightforward estimation procedures. Miao et al. (2023) consider a similar setting to (1) and (3), showing that the causal effect is identifiable if $\|\beta\|_{0}\leq(p-q)/2$. Their sparsity constraint is significantly stronger than ours, especially in cases where the number of exposures is large relative to the number of latent confounders. Miao et al. (2023) also develop a robust linear regression-based estimator for estimating $\beta$. In contrast to our estimator, their estimator is only consistent in the low-dimensional regime where $p$ is fixed and $\|\beta\|_{0}\leq p/2-q+1$. Furthermore, their estimator for $\beta$ is not sparse and, therefore, cannot be used for selecting treatments with non-zero effects.
@@
Our results also connect to recent literature on multiply robust causal identification (e.g., Sun et al., 2023), as we show identification in the union of many causal models. This contrasts with the rich literature on multiply robust estimators under the same causal model (e.g., Wang and Tchetgen Tchetgen, 2018) and improved doubly robust estimators that are consistent under multiple working models for two components of the likelihood (e.g., Han and Wang, 2013).
@@ 1.2 Outline of this paper
The rest of this article is organized as follows. In Section 2, we introduce the setup and background. In Section 3, we introduce our identification strategy using the synthetic instrument method. In Section 4, we present our estimation procedure and provide theoretical justifications. We also discuss extensions to non-linear outcome models. Simulation studies in Section 5 compare our proposal with several state-of-the-art methods in terms of their finite-sample performance. In Section 6, we apply our method to mouse obesity data. We conclude with a brief discussion in Section 7.
The proposed method is implemented in an R package available at https://github.com/dingketang/syntheticIV.
@@ 2.1 The model
We consider both low- and high-dimensional settings where $p$ may be smaller or larger than the sample size $n$. Assume $\dot{\beta}$ is the true value of $\beta$ in model (3). Without loss of generality, we assume all the variables in (1) and (3) are centered, $\text{Cov}(U)=I_{q}$, and $\mathbb{E}(g(U))=0$.
@@
Condition A1 is a regularity condition commonly assumed in the literature (e.g., Miao et al., 2023, Theorem 3). However, it may be relaxed in our setting. For example, if some treatment effects are not confounded after normalization, such that certain rows of $\text{Cov}^{-1}(X)\Lambda$ are zero, then although Condition A1 is violated, the proposed method can still be used to identify and estimate the treatment effects. See Section LABEL:sec:discussA1 of the supplementary material for more details.
Condition A2 has been discussed extensively in the factor model literature. One classical result is Proposition 1 below, which is a direct corollary of Anderson and Rubin (1956, Theorem 5.1).
@@
We note that the condition that $D$ is a diagonal matrix is a classical assumption in the factor analysis literature. However, Condition A2, and hence our algorithm, may still hold even if $D$ is not diagonal. For example, under the assumption that $D$ is sparse, the covariance structure $\text{Cov}(X)=D+\Lambda\Lambda^{\mathrm{\scriptscriptstyle T}}$ implies a sparse plus low-rank decomposition. This allows for the identification of the low-rank component $\Lambda\Lambda^{\mathrm{\scriptscriptstyle T}}$, as established in Chandrasekaran et al. (2011, Corollary 3), which leads to identifying $\Lambda$ up to a rotation. In another example, in the high-dimensional setting where $p\rightarrow\infty$, it is possible to identify $\Lambda^{*}\in\mathbb{R}^{p\times q}$, whose columns correspond to the top $q$ eigenvalues of $\text{Cov}(X)$. Under additional boundedness assumptions on the correlation matrix $D$ and the coefficient matrix $\Lambda$, one can show that there exists a matrix $O\in\mathbb{R}^{q\times q}$ such that the $\ell_{2}$ norm between each column of $\Lambda O$ and $\Lambda^{*}$ converges to zero as $p$ tends to infinity. See Fan et al. (2013, Proposition 2.2, Theorem 3.3), Bai (2003, Theorem 2), and Shen et al. (2016, Theorem 1) for more details.
@@ 2.2 Identifiability of the causal effect $\beta$
where $\Lambda_{j}$ is the $j$th element of $\Lambda\in\mathbb{R}^{p\times 1}$ and $\gamma=\mathbb{E}(Ug(U))$. Since there are three equations in (4) but four unknown parameters $\beta_{1},\beta_{2},\beta_{3},\gamma$, the causal parameters $\beta$ are not identifiable from these equations.
@@
    (Sparsity) $\|\dot{\beta}\|_{0}\leq p-q-1$, where $\dot{\beta}$ refers to the true value of $\beta$.
@@
Condition A3 is significantly less restrictive compared to similar assumptions in the existing literature used for causal effect identification in this context. For example, Miao et al. (2023) assumed $\|\dot{\beta}\|_{0}\leq(p-q)/2$.
@@ 2.3 Instrumental variable
The method of instrumental variables (IV) is a widely used approach for estimating causal relationships when unmeasured confounders exist between the exposure $X$ and the outcome $Y$. Suppose we have an exogenous variable $Z$. For simplicity, assume that the relationships among the random variables are linear and follow the structure equation models:
@@
For $Z$ to be a valid instrumental variable, the following assumptions are commonly made (e.g., Wang and Tchetgen Tchetgen, 2018): $\pi=0$ (exclusion restriction), $\alpha_{z}\neq 0$ (instrumental relevance), and $\text{Cov}(U,Z)=0$ (unconfoundedness). Under these assumptions, one can consistently estimate $\beta$ via a two-stage least squares estimator: first, obtain the predicted exposure $\widehat{\mathbb{E}}(X\mid Z)$ by linearly regressing $X$ on $Z$, and then regress $Y$ on $\widehat{\mathbb{E}}(X\mid Z)$ to obtain an estimate of $\beta$; here $\mathbb{E}(X\mid Z)$ refers to the conditional expectation of $X$ given $Z$.
@@ 3.1 A new identification approach via voting
To explain the identification strategy, it is helpful to consider a voting analogy; see also Zhou et al. (2014) and Guo et al. (2018) for similar approaches in different contexts. Suppose the analyst consults three experts, and expert $j$ hypothesizes that $\beta_{j}=0$. Based on this hypothesis, one can identify the other elements in $\beta$ using the approach described in Section 2.2. Specifically, for $j=1,2,3,$ let $\widetilde{\beta}^{(j)}$ (and $|\widetilde{\gamma}^{(j)}|$) solve (4) assuming $\beta_{j}=0$. Table 1 summarizes these solutions. Note that the hypotheses by experts 1 and 2 are both correct, so we have $\widetilde{\beta}^{(1)}=\widetilde{\beta}^{(2)}=\beta$ under Conditions A1A2. On the other hand, the hypothesis postulated by expert 3 is incorrect. Therefore, in general, $\widetilde{\beta}^{(3)}\neq\beta$. To decide among these three experts, we compare the solutions $\widetilde{\beta}^{(j)}$ and find their mode $\beta_{\text{mode}}=\mathop{\arg\max}\limits_{\beta\in\mathbb{R}^{3}}|\{j:\widetilde{\beta}^{(j)}=\beta\}|$; here $|\mathcal{S}|$ denotes the cardinality of a set $\mathcal{S}$. One can easily see from Table 1 that $\beta_{\text{mode}}=\dot{\beta}.$

### v3

@@ Abstract
In many observational studies, researchers are often interested in the effects of multiple exposures on a single outcome. Standard approaches for high-dimensional data, such as the Lasso, assume that the associations between the exposures and the outcome are sparse. However, these methods do not estimate causal effects in the presence of unmeasured confounding. In this paper, we consider an alternative approach that assumes the causal effects under consideration are sparse. We show that under sparse causation, causal effects are identifiable even with unmeasured confounding. Our proposal is built around a novel device called the synthetic instrument, which, in contrast to standard instrumental variables, can be constructed directly from the observed exposures. We demonstrate that, under the assumption of sparse causation, the problem of causal effect estimation can be formulated as an $\ell_{0}$-penalization problem and solved efficiently using off-the-shelf software. Simulations show that our approach outperforms state-of-the-art methods in both low- and high-dimensional settings. We further illustrate our method using a mouse obesity dataset.
@@
1Department of Mathematics and Statistics, University of Ottawa, Ottawa, Ontario, Canada
@@ 1 Introduction
Sparsity is a common assumption in the modern statistical learning literature, as it facilitates variable selection in models and enhances the interpretability of parameter estimates. For example, the Lasso (Tibshirani, 1996) assumes sparse associations between a single outcome and potentially high-dimensional predictors; in other words, only a small subset of predictors have nonzero associations with the outcome. Hastie et al. (2009) summarizes the philosophy behind such methods as the “bet on sparsity” principle: use a procedure that performs well in sparse settings, since no procedure performs well in dense ones. Methods like the Lasso perform well under sparse associations and, as a result, have gained significant popularity in recent decades.
Importantly, the “bet on sparsity” principle does not restrict the types of problems to which sparsity may apply. Beyond sparse associations, a growing body of literature emphasizes sparse causation, where only a fraction of exposures exert nonzero causal effects on the outcome (e.g., Spirtes and Glymour, 1991; Claassen et al., 2013; Wang et al., 2017; Miao et al., 2023; Zhou et al., 2024). This assumption is often more interpretable and plausible in real data applications. For example, suppose we are interested in the relationship between gene expression and a phenotype such as lung cancer. Biological evidence suggests that only a small proportion of genes may influence the risk of lung cancer (e.g., Kanwal et al., 2017). However, this does not imply sparse association, since unmeasured confounding may induce spurious correlations between many genes and the phenotype.
To illustrate, consider a linear structural model (Pearl, 2013) with a $p$-dimensional exposure vector $X=(X_{1},\ldots,X_{p})^{\mathrm{\scriptscriptstyle T}}$, an outcome $Y$, and a $q$-dimensional latent variable $U$:
@@
where $\Lambda\in\mathbb{R}^{p\times q}$, $\beta\in\mathbb{R}^{p}$, and $\gamma\in\mathbb{R}^{q}$ are coefficient vectors, and $\epsilon_{x}=(\epsilon_{1},\ldots,\epsilon_{p})^{\mathrm{\scriptscriptstyle T}}$, $\epsilon_{y}$, and $U$ are mutually uncorrelated. Under this model, spurious correlations induced by unmeasured confounding, given by $\text{Cov}(X)^{-1}\Lambda\gamma$, are typically dense. Consequently, the overall association between $X$ and $Y$ is dense, even when the causal effect $\beta$ itself is sparse.
Identification and estimation of the causal parameter $\beta$ are nontrivial due to the presence of unmeasured confounding by $U$. The contributions of this paper are twofold. First, under an additional plurality condition, we establish that the parameter $\beta$ in model (3) is identifiable if and only if $\|\beta\|_{0}<p-q$. This sparsity assumption is both necessary and sufficient, representing a significant improvement over conditions previously introduced in the literature; see Section 1.1 for details. Remarkably, in contrast to many other identification assumptions in causal inference, this assumption can be consistently tested from data.
Second, we develop a two-stage synthetic regularized regression approach for estimating $\beta$, with a first stage based on ordinary least squares and a second stage using $\ell_{0}$-penalized regression. The key technique behind our results is a novel device, which we term the synthetic instrument. Unlike standard instrumental variables, the synthetic instrument is constructed from a subset of exposures, enabling identification of causal effects without requiring exogenous variables. Our procedure enjoys Lasso-type theoretical guarantees in both low- and high-dimensional settings.
@@ 1.1 Related works
Our proposal is related to recent work on multivariate hidden confounding. Ćevid et al. (2020) and Guo et al. (2022) propose a spectral deconfounding method for estimating $\beta$ in a high-dimensional model. Their method assumes a dense confounding structure, which is feasible only in a high-dimensional regime where $p$ tends to infinity with the sample size and the magnitudes of spurious associations tend to zero. Bing et al. (2022) consider a more general setup than the one we study, in which they also allow the outcome to be multivariate. However, they aim to identify the projection of $\beta$ onto a related space rather than the causal parameter $\beta$ itself. Chandrasekaran et al. (2010) study a related problem under the assumption that $(X,Y)$ are normally distributed. Under this assumption, they not only identify the effect of $X$ on $Y$ but also recover the covariance among components of $X$ conditional on $U$.
@@
The estimation problem for $\beta$ can be framed within the context of causal inference with unmeasured confounding. Currently, the most popular approach in practice is the instrumental variable (IV) framework, which uses information from an exogenous variable known as an IV to identify causal effects (e.g., Angrist et al., 1996; Wang and Tchetgen Tchetgen, 2018; Pfister and Peters, 2022). Another approach that has gained attention recently is the proximal causal inference framework (Tchetgen Tchetgen et al., 2024), which uses information from ancillary variables, known as negative control exposures and outcomes, to remove bias due to unmeasured confounding. Compared with these frameworks, our approach does not rely on the collection of additional ancillary variables, which can be challenging in many practical settings. Instead, we rely on the availability of multiple exposures and the sparsity assumption for identification and estimation.
@@
Recently, a strand of literature has sought to identify the causal effects of multiple exposures. Wang and Blei (2019) popularized this setting by proposing the so-called deconfounder method, which first obtains an estimate $\widehat{U}$ of the unmeasured confounder and then adjusts for $\widehat{U}$ using standard regression methods. However, it has been pointed out that in this setting, without further assumptions, the causal effect $\beta$ is not identifiable (D’Amour, 2019; Ogburn et al., 2020). Kong et al. (2022) show that under model (1) and a binary choice model for the outcome with a non-probit link, the causal effects are identifiable. Their identification results, however, apply only to binary outcomes and do not lead to straightforward estimation procedures. Miao et al. (2023) consider a similar setting to (1) and (3), showing that the causal effect is identifiable if $\|\beta\|_{0}\leq(p-q)/2$. Their sparsity constraint is significantly stronger than ours, especially when the number of exposures is large relative to the number of latent confounders. Miao et al. (2023) also develop a robust linear regression-based estimator for $\beta$. In contrast to our estimator, their estimator is consistent only in the low-dimensional regime where $p$ is fixed and $\|\beta\|_{0}\leq p/2-q+1$. Furthermore, their estimator for $\beta$ is not sparse and therefore cannot be used for selecting treatments with nonzero effects.
@@
Our results also connect to recent literature on multiply robust causal identification (e.g., Sun et al., 2023), as we show identification in the union of many causal models. This contrasts with the extensive literature on multiply robust estimators under the same causal model (e.g., Wang and Tchetgen Tchetgen, 2018) and on improved doubly robust estimators that are consistent under multiple working models for two components of the likelihood (e.g., Han and Wang, 2013).
@@ 1.2 Outline of this paper
The rest of this article is organized as follows. In Section 2, we introduce the setup and background. In Section 3, we describe our identification strategy using the synthetic instrument method. In Section 4, we present our estimation procedure and provide theoretical justifications. We also discuss extensions to nonlinear outcome models. Simulation studies in Section 5 compare our proposal with several state-of-the-art methods in finite-sample performance. In Section 6, we apply our method to mouse obesity data. We conclude with a brief discussion in Section 7.
The proposed method is implemented in an R package, available at https://github.com/dingketang/syntheticIV.
@@ 2.1 The model
We consider both low- and high-dimensional settings, where $p$ may be smaller or larger than the sample size $n$. Let $\dot{\beta}$ denote the true value of $\beta$ in model (3). Without loss of generality, we assume all the variables in (1) and (3) are centered, $\text{Cov}(U)=I_{q}$, and $\mathbb{E}(g(U))=0$.
@@
Condition A1 is a regularity condition commonly assumed in the literature (e.g., Miao et al., 2023, Theorem 3). However, it may be relaxed in our setting. For example, if certain treatment effects are unconfounded after normalization, so that specific rows of $\text{Cov}^{-1}(X)\Lambda$ are zero, then even though Condition A1 is violated, our proposed method can still be used to identify and estimate the treatment effects. See Sections LABEL:sec:discussA1 and LABEL:sec:weakivandidentification of the supplementary material for further details on how the algorithm identifies treatment effects under relaxed versions of A1.
Condition A2 has been discussed extensively in the factor model literature. One classical result is Proposition 1, which is a direct corollary of Anderson and Rubin (1956, Theorem 5.1).
@@
We note that the condition that $D$ be a diagonal matrix is a classical assumption in the factor analysis literature. However, Condition A2, and hence our algorithm, may still hold even if $D$ is not diagonal. For example, under the assumption that $D$ is sparse, the covariance structure $\text{Cov}(X)=D+\Lambda\Lambda^{\mathrm{\scriptscriptstyle T}}$ implies a sparse plus low-rank decomposition. This allows for the identification of the low-rank component $\Lambda\Lambda^{\mathrm{\scriptscriptstyle T}}$, as established in Chandrasekaran et al. (2011, Corollary 3), which leads to identifying $\Lambda$ up to a rotation. In another example, in the high-dimensional setting where $p\rightarrow\infty$, it is possible to identify $\Lambda^{*}\in\mathbb{R}^{p\times q}$, whose columns correspond to the top $q$ eigenvalues of $\text{Cov}(X)$. Under additional boundedness assumptions on the correlation matrix $D$ and the coefficient matrix $\Lambda$, one can show that there exists a matrix $O\in\mathbb{R}^{q\times q}$ such that the $\ell_{2}$-norm between each column of $\Lambda O$ and $\Lambda^{*}$ converges to zero as $p$ tends to infinity. See Fan et al. (2013, Proposition 2.2, Theorem 3.3), Bai (2003, Theorem 2), and Shen et al. (2016, Theorem 1) for more details.
@@ 2.2 Identifiability of the causal effect $\beta$
where $\Lambda_{j}$ is the $j$th element of $\Lambda\in\mathbb{R}^{p\times 1}$ and $\gamma=\mathbb{E}(Ug(U))$. Since there are three equations in (4) but four unknown parameters, $\beta_{1},\beta_{2},\beta_{3},\gamma$, the causal parameters $\beta$ are not identifiable from these equations.
@@
    (Sparsity) $\|\dot{\beta}\|_{0}\leq p-q-1$, where $\dot{\beta}$ denotes the true value of $\beta$.
@@
Condition A3 is significantly less restrictive than similar assumptions in the existing literature used for causal effect identification in this context. For example, Miao et al. (2023) assumed $\|\dot{\beta}\|_{0}\leq(p-q)/2$.
@@ 2.3 Instrumental variable
The method of instrumental variables is a widely used approach for estimating causal relationships when unmeasured confounders exist between the exposure $X$ and the outcome $Y$. Suppose we have an exogenous variable $Z$. For simplicity, assume that the relationships among the random variables are linear and follow the structural equation models:
@@
For $Z$ to be a valid instrumental variable, the following assumptions are commonly made (e.g., Wang and Tchetgen Tchetgen, 2018): $\pi=0$ (exclusion restriction), $\alpha_{z}\neq 0$ (instrumental relevance), and $\text{Cov}(U,Z)=0$ (unconfoundedness). Under these assumptions, one can consistently estimate $\beta$ via a two-stage least squares estimator: first, obtain the predicted exposure $\widehat{\mathbb{E}}(X\mid Z)$ by linearly regressing $X$ on $Z$, and then regress $Y$ on $\widehat{\mathbb{E}}(X\mid Z)$ to obtain an estimate of $\beta$. Here, $\mathbb{E}(X\mid Z)$ refers to the conditional expectation of $X$ given $Z$, and $\widehat{\mathbb{E}}(X\mid Z)$ refers to its estimator obtained through linear regression.
@@ 3.1 A new identification approach via voting
To explain the identification strategy, it is helpful to consider a voting analogy; see also Zhou et al. (2014) and Guo et al. (2018) for similar approaches in different contexts. Suppose the analyst consults three experts, and expert $j$ hypothesizes that $\beta_{j}=0$. Based on this hypothesis, one can identify the other elements in $\beta$ using the approach described in Section 2.2. Specifically, for $j=1,2,3$, let $\widetilde{\beta}^{(j)}$ (and $|\widetilde{\gamma}^{(j)}|$) solve (4) assuming $\beta_{j}=0$. Table 1 summarizes these solutions. Note that the hypotheses by experts 1 and 2 are both correct, so we have $\widetilde{\beta}^{(1)}=\widetilde{\beta}^{(2)}=\beta$ under Conditions A1A2. On the other hand, the hypothesis postulated by expert 3 is incorrect. Therefore, in general, $\widetilde{\beta}^{(3)}\neq\beta$. To decide among these three experts, we compare the solutions $\widetilde{\beta}^{(j)}$ and find their mode, defined as $\beta_{\text{mode}}=\mathop{\arg\max}\limits_{\beta\in\mathbb{R}^{3}}|\{j:\widetilde{\beta}^{(j)}=\beta\}|,$ where $|\mathcal{S}|$ denotes the cardinality of a set $\mathcal{S}$. One can easily see from Table 1 that $\beta_{\text{mode}}=\dot{\beta}$.